# Elastic Metastability: Thermally Activated Snap-through Transitions in Nanostructures

### v1

@@ Abstract
In parallel, unbiased MD simulations are conducted to record the escape times for transitions from the inverted state to the natural state at 300 K, 325 K, 350 K, and 375 K. At each temperature, the transition rate constant $k$ is obtained as the inverse of the mean escape time $\tau_{e}$ from 100 independent trajectories (see the Supplemental Material). To analyze the temperature dependence of $k_{\mathrm{TST}}$, we recall that $w_{0}$ is computed as a linear combination of the midline C atom coordinates. This construction renders $\langle{Z_{w_{0}}}^{\frac{1}{2}}\rangle_{w_{0}={w_{0}}^{\ddagger}}$ effectively constant, leaving the temperature dependence of $k_{\mathrm{TST}}$ governed by $\beta$ in both the prefactor and the Boltzmann factor of Eq. (4). Motivated by this observation, we plot $\ln{(k\,T^{-\frac{1}{2}})}$ against $1/T$ in Fig. 5(a). The fit to the data exhibits good linearity, indicating that $k$ follows the temperature dependence described by Eq. (4). Moreover, the slope of the fit gives $\Delta F^{\ddagger}$ = 51.228 meV, in good agreement with the $\Delta F^{\ddagger}$ values previously calculated from WT-MetaD (average relative error $5\%$). These findings demonstrate that the temperature dependence of the rate constant $k$ is accurately captured by the generalized TST, and corroborate the self-consistency of our computational results. In practice, the actual rate constant $k$ generally deviates from the TST prediction $k_{\mathrm{TST}}$ by a multiplicative factor, the transmission coefficient $\kappa$, which accounts for recrossing effects. However, the consistent temperature dependence exhibited by $k$ and $k_{\mathrm{TST}}$ implies that, for our system, $\kappa$ remains approximately constant over at least a moderate range of temperature.
@@
While generalized TST alone does not directly yield the absolute value of the rate constant $k$, Eqs. (3) and (4) still allow us to predict the kinetic behavior of the GNR system under various conditions by comparing relative changes in $\Delta F^{\ddagger}$ and temperature. In this context, we estimate the mean escape time $\tau_{e}$ for the forward transition at various temperatures, with the GNR constrained at $\alpha=24^{\circ}$ (Fig. 5(b)). A temperature variation of $\Delta T=20$ K within the studied range can result in orders-of-magnitude changes in $\tau_{e}$ and hence in the rate constant $k$. In addition, the backward transition is associated with $\Delta F_{\text{back}}^{\ddagger}=2.31$ eV at 300 K, corresponding to a mean escape time of $\sim 10^{23}$ years, which indicates that the natural state is effectively stable under ambient conditions. This analysis highlights that temperature control, along with adjusting the free energy via geometric constraints (Fig. 3(b)), can offer an effective strategy for tuning the transition rate, its temperature sensitivity, and the transition direction of the thermally activated snap-through in the GNR system.
In conclusion, we have shown that snap-through transitions in geometrically constrained GNRs can occur as well-defined thermally activated processes. By biasing the order parameter in WT-MetaD simulations, we have precisely characterized the free energy landscape, transition pathways, and the Landau free energy of the bistable GNR system. The temperature dependence of the transition rate constant is accurately captured by reaction rate theory. Our findings can be generalized to other bistable elastic nanostructures coupled to a thermalizing environment. With appropriately designed size, temperature, and structural energetics, such systems can exhibit morphological transitions via thermal activation at controllable transition rates. This study lays the foundation for understanding elastic metastability and opens up potential avenues for designing nanoscale thermal switches, thermal actuators, and other components in temperature-responsive devices.
This work has been supported by the National Natural Science Foundation of China under Grants No. 21673119 and No. 21973046.

### v3

@@ Results
In parallel, unbiased MD simulations are conducted to record the escape times for transitions from the inverted state to the natural state at 300 K, 325 K, 350 K, and 375 K, respectively. At each temperature, the transition rate constant $k$ is obtained as the inverse of the mean escape time $\tau_{e}$ from 100 independent trajectories (see the Supporting Information). To analyze the temperature dependence of $k_{\mathrm{TST}}$, we recall that $w_{0}$ is computed as a linear combination of the midline C atom coordinates, leaving $\langle{Z_{w_{0}}}^{\frac{1}{2}}\rangle_{w_{0}={w_{0}}^{\ddagger}}$ constant, and the temperature dependence of $k_{\mathrm{TST}}$ lies in the prefactor $T^{\frac{1}{2}}$ and the Boltzmann factor of Eq. (4). Motivated by this observation, we plot $\ln{(k\,T^{-\frac{1}{2}})}$ against $1/T$ in Fig. 5(a), where a linear trend indicates that $k$ follows the temperature dependence described by Eq. (4). Moreover, the slope of the linear fit gives $\Delta F^{\ddagger}$ = 51.228 meV, which is in good agreement with the $\Delta F^{\ddagger}$ values previously calculated from WT-MetaD (average relative error $5\%$). These findings demonstrate that the temperature dependence of the rate constant $k$ is accurately captured by generalized TST, and corroborate the self-consistency of our computational results. In practice, the rate constant $k$ generally deviates from the TST prediction $k_{\mathrm{TST}}$ by a multiplicative factor, the transmission coefficient $\kappa$, which accounts for recrossing effects. However, the consistent temperature dependence exhibited by $k$ and $k_{\mathrm{TST}}$ here implies that, for our system, $\kappa$ remains approximately constant over at least a moderate range of temperature.
While generalized TST alone does not directly yield the absolute value of the rate constant $k$, Eq. (3) still allows us to predict the kinetics of the GNR system under various conditions by comparing relative changes in $F(w_{0})$ and temperature. To illustrate this, the geometric constraint is adjusted from $\alpha=29^{\circ}$ to $24^{\circ}$, and the updated $F(w_{0})$ is calculated with WT-MetaD at room temperature. Based on the difference of $F(w_{0})$ in the two settings, Eq. (3) predicts $k=66.68$ ms-1 at $\alpha=24^{\circ}$ and $T$ = 300 K. Meanwhile, we compute the rate constant by the accelerated MD of hyperdynamics [voter1997method, voter1997hyperdynamics] under the same conditions, obtaining $k=77.69\pm 21.54$ ms-1 (see the Supporting Information). The prediction of Eq. (3) falls within the 95% confidence interval of the hyperdynamics computation result. Thus, the predictive performance of generalized TST is robust across variations in the free energy landscape that are realized by changing the geometric confinement.
@@
Fig. 5(b) presents the evolution of the rate constant $k$ computed by Eq. (3) based on the WT-MetaD free energy calculations at $\alpha=24^{\circ}$ and various temperatures. A temperature variation of $\Delta T=25$ K can result in up to orders-of-magnitude changes in $k$. In addition, the backward transition is associated with $\Delta F_{\text{back}}^{\ddagger}=2.31$ eV at 300 K, corresponding to a mean escape time of $\sim 10^{23}$ years, which indicates that the natural state is effectively stable under this condition. This construction can be exploited to control the transition direction and is particularly useful in the application of a thermal switch or actuator. Our analysis highlights that temperature control, along with the adjustment of activation free energy through geometric confinements (as illustrated in Fig. 5(b)), can offer an effective strategy for tuning the transition rate, its temperature sensitivity, and the transition direction of the activated snap-through in the GNR system.
Discussion
The effects of thermal fluctuations on 2D materials are ubiquitous, but thermally activated morphological transitions in these systems are uncommon under ambient conditions. This is because realizing a well-defined activated process typically requires an activation free energy of a certain order of magnitude to produce observable transitions on accessible timescales. The GNR system discussed in this work exhibits a free energy landscape that is tunable through geometric confinement, thereby providing a prototype for realizing and regulating thermally activated morphological transitions in 2D materials. Moreover, recent years have seen a surge of interest in bistable elastic structures that display snap-through behavior [gomez2017critical, gomez2018pull, pandey2014dynamics, sano2017slip, sano2018snap, radisson2023elastic, radisson2023dynamic, wang2024transient]. When scaled down to the nanoscale, such structures may offer metastable free energy landscapes controlled by external constraints, suggesting a promising route for engineering thermally activated mechanical transitions at the nanoscale.
In this work, we introduce enhanced sampling methods, including WT-MetaD and hyperdynamics into the study of nanoscale elastic systems exhibiting metastability. WT-MetaD enables us to reconstruct the complex free energy landscape and to reveal a pair of asymmetric transition pathways of the geometrically constrained GNR system. The resultant Landau free energy profile is compatible with the application of generalized transition state theory, which allows us to accurately capture the kinetics of the system over a range of temperature and geometric constraints. This methodology can be readily generalized to other bistable elastic nanostructures coupled to a thermalizing environment. For more complex structures, even in the absence of prior knowledge about the optimal collective variables, data-driven collective variable construction methods can be integrated with WT-MetaD to recover nontrivial free energy landscapes [wang2020machine, wang2019past, mehdi2024enhanced, zhu2025enhanced]. Our findings offer a theoretical framework for the study of elastic metastability, advance the understanding of thermalized nanomechanical systems, and provide design strategies for tunable nanoscale thermal switches and thermal actuators.
Materials and Methods
All simulations were performed with the LAMMPS [plimpton1995fast] package. The atomic interactions of the studied system were described by the REBO-II potential [brenner2002second], which is one of the most commonly used empirical potentials for solid carbon and hydrocarbon systems and has been shown to reliably capture the elasticity [lu2009elastic] and thermomechanical behaviors [gao2014thermomechanics] of graphene. The outermost seven rows of atoms at each end of the GNR were rigidly held to implement the geometric constraint, and the other atoms evolved according to the velocity-Verlet algorithm with a time step of 1.0 fs. A Langevin thermostat was used to provide the thermalizing environment at specified temperatures with a damping coefficient of 1.0 ps-1 [savin2022friction, li2019influence, savin2023critical]. The initial configurations of unbiased simulations were obtained from sampling the equilibrium ensemble of the inverted state or the natural state. WT-MetaD and hyperdynamics simulations were carried out with the COLVARS module [fiorin2013using]. The WT-MetaD parameters were optimized when the system boundaries were constrained at each angle $\alpha$. The order parameter $w_{0}$ was computed as the transverse displacement of the center of mass of the C atoms in the longitudinal middle row of the GNR. The collective variable $\Delta w_{\text{qrt}}$ was defined as $\Delta w_{\text{qrt}}=w(L/4)-w(-L/4)$, where $w(L/4)$ and w(L/4)w(-L/4) denote the average transverse displacements of the C atom rows located at the quarter-length positions measured from each end of the GNR. Further methodological and computational details can be found in the Supporting Information.
Acknowledgements.
R.Z., Y.Z. and C.L. gratefully acknowledge support from the National Natural Science Foundation of China under Grant No. 21973046. Y.W. gratefully acknowledges support from the startup funds provided by Case Western Reserve University.